Hi
from a lighthouse 70m above sea level a ship 1.2km out to sea is observed. What is the angle of
depression from the lighthouse to the ship.
thanks

Found 2 solutions by rothauserc, MathTherapy:
Answer by rothauserc(4718)

1.2km = 1200m
:
tan theta = 1200 /70 = 17.1429
:
tan^-1 17.1429 = 86.6615 degrees
:
***************************************************
angle of depression = 90 - 86.6615 = 3.3385 degrees
***************************************************
:

Answer by MathTherapy(10553)


Hi
from a lighthouse 70m above sea level a ship 1.2km out to sea is observed. What is the angle of
depression from the lighthouse to the ship.
thanks
1.2km = 1,200 m
Angle of DEPRESSION from lighthouse = angle of ELEVATION from ship
We then get: